Please help :)
given: u = sqrt(x^2+y^2)
solve: [(du/dx)/(du/dy)]

Answer by stanbon(75887)

given: u = sqrt(x^2+y^2)
solve: [(du/dx)/(du/dy)]
-------
du/dx = (1/2)(2x)/sqrt(x^2+y^2) = x/sqrt(x^2+y^2)
-------
du/dy = (1/2)(2y)/sqrt(x^2+y^2) = y/sqrt(x^2+y^2)
----
Ans: x/y
=====
Cheers,
Stan H.
==========